OpenStudy (anonymous):

2a^2 + 30a + 100 Is it prime when factored?

OpenStudy (anonymous):

i got prime

OpenStudy (anonymous):

when i did it

OpenStudy (anonymous):

so you factored to this? \[2a^2 + 30a + 100 = (2a +10)(a+10) = 2(a+5)(a+10)\]

OpenStudy (anonymous):

well isnt that wrong?

OpenStudy (anonymous):

im not sure quite what you mean by prime in this case

OpenStudy (anonymous):

well, can you factor this equation?

OpenStudy (anonymous):

i cant seem to

OpenStudy (anonymous):

yes i did so above :D

OpenStudy (anonymous):

i'll help

OpenStudy (anonymous):

because there are no factors of ten that add to 30

OpenStudy (anonymous):

i mean 100 not 10

OpenStudy (anonymous):

ah i see your problem, here we go: we have 2a^2 and not just a^2 , so when we factor it will effect what numbers we are looking for. we are going to have (2a + (something) )( a + (something else) ) so you're right the numbers we are looking for will multiply to 100 but we need to think about the 2a term when we think how they add to get 30 10*10 = 100 and 2*10 + 10 = 30 so it becomes (2a+10)(a+10) you can check by multiplying out

OpenStudy (anonymous):

ok thanks! :D:D

OpenStudy (anonymous):

:) happy to help

OpenStudy (anonymous):

wait hold on

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

sorry, but why do you do the 2 * 10 + 10 = 30 part. Where do you get it from?

OpenStudy (anonymous):

its from how we multiply out the brackets, because factorising is just the reverse of expanding: lets pretend we dont know about the 10's we are looking for two numbers C and D first we see that one of the a's in the brackets has to be multiplied by 2: 2a^2 + 30a + 100 = (2a + C)(a + D) lets expand that right hand side: (2a + C)(a + D) = 2a(a+D) + C(a+D) = 2a^2 + 2aD + aC + CD now we know 2a^2 + 30a + 100 = 2a^2 + 2aD + aC + CD so CD = 100 and 2aD + aC = 30a so our numbers multiply to get 100 and if we double one of them and add the other we get 30 im really sorry if this has confused you more

OpenStudy (anonymous):

hmmm...

OpenStudy (anonymous):

i sort of get this.

OpenStudy (anonymous):

prime.

OpenStudy (anonymous):